a step index fiber with large core diameter compared with the wavelength of the transmitted light has an acceptance angle in air of 22 and a relative refractive index difference of 3%.Estimate the numerical aperture and the critical angle at the core-cladding interface for the fiber
Answer #1

The numerical aperture can be calculated using the formula NA = Sin(22) where 22 is abgle of acceptance, so the NA comes out to be -0.008513092

the cirtical angle can be calculated using the SinC= 1/relative difference == C=sin-1(1/(3/100))

C=sin-1(100/3)